OpenStudy (anonymous):

Factor completely: 10x2 + 39x + 14 A. (10x + 7)(x + 2) B. (5x + 2)(2x + 7) C. (5x + 7)(2x + 2) D. (10x + 2)(x + 7)

OpenStudy (inkyvoyd):

Well, for these problems, you can actually just expand A,B,C,and D.

mathslover (mathslover):

@Senior012 inkyvoyd told u a great hint : \[\huge{10x^2+39x+140}\] solve it through quadratic equation : \(\huge{ax^2+bx+c=0}\) -----------------GENERAL FORM OF QUADRATIC EQUATION \[\huge{x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}}\] a = 10 , b = 39 , c = 140

mathslover (mathslover):

solve it by putting the values of a , b and c that is \[\huge{x=\frac{-39\pm\sqrt{(-39)^2-4(10)(140)}}{2(10)}}\] u can solve this ..

mathslover (mathslover):

may i know @Senior012 did u understand this ?

OpenStudy (anonymous):

NO /:

mathslover (mathslover):

so where r u having problem ?

OpenStudy (anonymous):

The right answer is B you have to break the middle term i.e 39x Into two values in such a way that the two values when added give 39 and when multiple give 140(the product of 10x2 and 14). The two values are: 35 and 4 place them in place of 39x and solve it :-) 10x2+39x+14 10x2+35x+4x+14 Take out the common from first two and then from the other two 5x(2x+7)+2(2x+7) (5x+2)(2x+7)